Suppose that the British economy produces two goods: laptops and books. The quantity produced and the prices of these items for 2015 and 2016 are shown in the table below: Year Quantities produced Price ($) 2015 Laptops =50 Books =1,000 Laptops =200 Books = ? 2016 Laptops =100 Books =? Laptops =150 Books =10Suppose that the British economy produces two goods: laptops and books. The quantity produced and the prices of these items f
Answer #1

Given nominal GDP of Britain in 2015 = $ 15,000

Note nominal GDP is equal to real GDP.

NGDP2015 = 50 x 200 + 1,000 x x

15.000 = 10,000+ 1.000 x x

1.000 x x = 15,000 – 10,000

5,000 1.000

= x = $5

Book price in year 2015 = $ 5

B. In 2016 Nominal GDP increased by 10%.

NGDP of 2016 = $ 16,500

NGDP2016 = 100 x 150+ x x 10

16.500 = 100 x 150 + 2 x 10

10x = 16.500 – 15.000

BC= 1,500 10

= x= 150

C. Real gdp of 2015 = Nominal GDP of 2015 = $ 15,000

Real GDP of 2016 = 100 × 200 + 150 × 5 = $ 20,750

Growth rate = (20,750/15,000) - 1 = 0.3833

Growth rate = 38.33%